# Reinforcement Learning from Human Feedback without Reward Inference: Model-Free Algorithm and Instance-Dependent Analysis

### v1

@@ 1 Introduction
Reinforcement learning (RL), with a wide range of applications in gaming AIs (Knox & Stone, 2008; MacGlashan et al., 2017; Warnell et al., 2018), recommendation systems (Yang et al., 2022; Zeng et al., 2016; Kohli et al., 2013), autonomous driving (Wei et al., 2023; Schwarting et al., 2018; Kiran et al., 2021) , and large language model (LLM) training (Wu et al., 2021; Nakano et al., 2021; Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020), has achieved tremendous success in the past decade. A typical reinforcement learning problem involves an agent and an environment, where at each step, the agent observes the state, takes a certain action, and then receives a reward signal. The state of the environment then transits to another state, and this process continues. However, most RL advances remain in the simulator environment where the data acquisition process heavily depends on the crafted reward signal, which limits RL from more realistic applications such as LLM, as defining a universal reward is generally difficult. In recent years, using human feedback as reward signals to train and fine-tune LLMs has delivered significant empirical successes for AI alignment problems and produced dialog AIs such as the ChatGPT (Ouyang et al., 2022). This paradigm where the reward of the state and actions is inferred from real human preferences, instead of being handcrafted, is referred to as Reinforcement Learning from Human Feedback (RLHF). A typical RLHF algorithm on LLMs involves three steps: (i) pre-train a network with supervised learning, (ii) infer a reward model from human feedback, in the form of comparisons or rankings among trajectories (responses), and (iii) use classic RL algorithm to fine-tune the pre-trained model. An accurate reward model that aligns with human preferences is the key to the superiority of RLHF.
@@ 2 Preliminaries
where $u:\mathbb{R}\to[0,1]$ is referred to as the link function (Bengs et al., 2021) which characterizes the structure of preference models. Other link functions, such as linear function, probit function, cloglog function, and cauchit function, have also been well-studied in dueling bandits (Ailon et al., 2014) and generalized linear models (Razzaghi, 2013; McCullagh, 2019), but not RLHF. In this paper, we use a 00-$1$ link function that indicates the favored set with higher reward, i.e.,
@@
Problem Formulation: Our goal is to design a learning algorithm to interact with the MDP and learn the optimal policy $\pi^{*}$ from the human feedback as quickly as possible. A learning algorithm $\mathsf{Alg}$ consists of (i) a sampling rule which decides which policy to choose at each episode and whether to query the human agent, (ii) a stopping rule which decides a stopping time when the learner wishes to output a learned policy, and (iii) a decision rule which decides which policy $\hat{\pi}$ to output. We call an algorithm $\delta$-PAC if it outputs an optimal policy with probability at least $1-\delta$. Our goal is to design such an algorithm to minimize sample complexity $K$:
@@ 4 Generalization to Discounted MDPs
$K=\tilde{\mathcal{O}}\left(\frac{SA^{3}M\log(\frac{1}{\delta})\log^{3}(\frac{1}{\varepsilon})}{(1-\gamma)^{3}(\overline{\Delta}_{\min}^{M})^{2}\max_{\pi}\min_{s^{\prime},h}p^{\pi}_{h}(s|s^{\prime})^{2}}\right),$
where $\overline{\Delta}_{\min}^{M}$ is min probability gap and $p^{\pi}_{h}(s|s^{\prime})$ is the visitation of $s$ after $h$ steps starting from $s^{\prime}$.
@@ Appendix A Related Work
Reinforcement learning, with a wide range of applications in gaming AIs (Knox & Stone, 2008; MacGlashan et al., 2017; Warnell et al., 2018) , recommendation systems (Yang et al., 2022; Zeng et al., 2016; Kohli et al., 2013) , autonomous driving (Wei et al., 2023; Schwarting et al., 2018; Kiran et al., 2021) , and large language model (LLM) training (Ziegler et al., 2019), has achieved tremendous success in the past decade. A typical reinforcement learning problem involves an agent and an environment, where at each step, the agent observes the state, takes a certain action, and then receives a reward signal. The state of the environment then transits to another state, and this process continues. In this section, we review works on RL that is relevant to our paper.
@@
Adaptive Stopping Design in RL: in best arm identification, usually, the key to delivering a promising and close to optimal sample complexity performance is to design the stopping criteria which stops the sampling rule as soon as the best action is identifiable. Before “model-based” methods such as TAS (Garivier & Kaufmann, 2016) which uses complicated Chernoff statistics to control the stopping time, there were many “confidence-based” algorithms (Kalyanakrishnan et al., 2012; Kaufmann & Kalyanakrishnan, 2013; Gabillon et al., 2012) focused on constructing high-probability confidence intervals. The stopping time is then when one confidence interval is disjoint from and greater than all the rest, which inspires our design. A similar designing principle is also adopted for bandit algorithms with multiple objectives, e.g., regret minimization while identifying the optimal action (Degenne et al., 2019; Zhang & Ying, 2024). Within this approach, the algorithms can identify the exact optimal policy, while also adapts the sample complexity to the structure of the problem, involving constants such as the value function gap in the theoretical performance bounds.
@@
Reinforcement Learning from Human Feedback: In recent years, the development of LLMs (Wu et al., 2021; Nakano et al., 2021; Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020) has motivated the study of RLHF (Christiano et al., 2017; Kaufmann et al., 2023), which is a generalization of dueling bandits in the MDP setting. Xu et al. (2020) studies the RLHF problem in standard tabular MDPs by reducing the RL problem into bandits of each state. Novoseller et al. (2020) generalizes posterior sampling to preference-based RL and Saha et al. (2023) studies RLHF under trajectory feature and preferences. For MDPs with function approximation, most existing work (Saha et al., 2023; Zhan et al., 2023b; a; Chen et al., 2022; Kong & Yang, 2022) are based on the Bradley-Terry-Luce model and MLE to characterize the human preferences and transform it into rewards. Specifically, in (Wang et al., 2023), the authors build a preference to reward interface that uses confidence balls to decide whether to query human feedback. Zhu et al. (2023) generalizes reward inference from pair-wise comparison to multiple preference rankings. (Kausik et al., 2024) studies RLHF with partially observed rewards and states. All aforementioned RLHF algorithms are value-based, and Du et al. (2024) studies policy gradient with human feedback. To the best of our knowledge, almost all works in literature adopt the pipeline that infers the reward model and then performs classic RL algorithms. Attempts to bypass reward modeling have also been studied. In (Rafailov et al., 2024), the authors proposed the DPO method which optimizes the policy directly from the preference over trajectories. However, DPO is purely empirical and still assumes the Bradley-Terry model to implicitly infer the reward. Moreover, DPO requires the knowledge of a reference policy and thus is only suitable in the fine-tuning phase. Attempts to understand the theoretical performance of DPO and its generalizations are made in (Azar et al., 2024), but the authors showed the existence of optima of the loss function, which does not provide any policy optimality and sample complexity guarantees.
@@ C.1 $\mathsf{BSAD}$ for Discounted MDPs
Assume the Condorcet winner exists for all state $s$ and trajectory length $h\leq H$. We let $\overline{\Delta}_{h}^{M}(s,a)$ to be the probability gap for action $a$ and trajectories of length $h$ compared to the Condorcet winner of that state $s$, and $p^{\pi}_{h}(s|s^{\prime})$ is the visitation probability of $s$ after $h$ steps starting from state $s^{\prime}$ with policy $\pi$. Both definitions are analogs to the definitions in episodic MDPs.
@@
$K=\tilde{\mathcal{O}}\left(\frac{SA^{3}M\log(\frac{1}{\delta})\log^{3}(\frac{1}{\varepsilon})}{(1-\gamma)^{3}\min_{h,s,a}\overline{\Delta}_{h}^{M}(s,a)^{2}\max_{\pi}\min_{s^{\prime},h}p^{\pi}_{h}(s|s^{\prime})^{2}}\right).$

### v2

@@ 1 Introduction
Reinforcement learning (RL), with a wide range of applications in gaming AIs (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Warnell et al., 2018), recommendation systems (Yang et al., 2024; Zeng et al., 2016; Kohli et al., 2013), autonomous driving (Wei et al., 2024; Schwarting et al., 2018; Kiran et al., 2022) , and large language model (LLM) training (Wu et al., 2021; Nakano et al., 2021; Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020), has achieved tremendous success in the past decade. A typical reinforcement learning problem involves an agent and an environment, where at each step, the agent observes the state, takes a certain action, and then receives a reward signal. The state of the environment then transits to another state, and this process continues. However, most RL advances remain in the simulator environment where the data acquisition process heavily depends on the crafted reward signal, which limits RL from more realistic applications such as LLM, as defining a universal reward is generally difficult. In recent years, using human feedback as reward signals to train and fine-tune LLMs has delivered significant empirical successes for AI alignment problems and produced dialog AIs such as the ChatGPT (Ouyang et al., 2022). This paradigm where the reward of the state and actions is inferred from real human preferences, instead of being handcrafted, is referred to as Reinforcement Learning from Human Feedback (RLHF). A typical RLHF algorithm on LLMs involves three steps: (i) pre-train a network with supervised learning, (ii) infer a reward model from human feedback, in the form of comparisons or rankings among trajectories (responses), and (iii) use classic RL algorithm to fine-tune the pre-trained model. An accurate reward model that aligns with human preferences is the key to the superiority of RLHF.
@@ 2 Preliminaries
where $u:\mathbb{R}\to[0,1]$ is referred to as the link function (Bengs et al., 2021) which characterizes the structure of preference models. Other link functions, such as linear function, probit function, cloglog function, and cauchit function, have also been well-studied in dueling bandits (Ailon et al., 2014) and generalized linear models (Razzaghi, 2013; McCulloch, 2000), but not RLHF. In this paper, we use a 00-$1$ link function that indicates the favored set with higher reward, i.e.,
@@
Problem Formulation: Our goal is to design a learning algorithm to interact with the MDP and learn the optimal policy $\pi^{*}$ from the human feedback as quickly as possible. A learning algorithm $\mathsf{Alg}$ consists of (i) a sampling rule which decides which policy to choose at each episode and whether to query the human agent, (ii) a stopping rule which decides a stopping time when the learner wishes to output an learned policy, and (iii) a decision rule which decides which policy $\hat{\pi}$ to output. We call an algorithm $\delta$-PAC if it outputs an optimal policy with probability at least $1-\delta$. Our goal is to design such an algorithm to minimize sample complexity $K$:
@@ 4 Generalization to Discounted MDPs
$K=\tilde{\mathcal{O}}\left(\frac{SA^{3}M\log(\frac{1}{\delta})\log^{3}(\frac{1}{\varepsilon})}{(1-\gamma)^{3}\min_{h,s,a}\overline{\Delta}_{h}^{M}(s,a)^{2}\max_{\pi}\min_{s^{\prime}}p^{\pi}_{h}(s|s^{\prime})^{2}}\right),$
where $\overline{\Delta}_{h}^{M}(s,a)$ to be the probability gap for action $a$ and trajectories of length $H-h$ compared to the Condorcet winner of that state $s$, and $p^{\pi}_{h}(s|s^{\prime})$ is the visitation probability of $s$ after $h$ steps starting from state $s^{\prime}$ with policy $\pi$. Both definitions are analog to the definitions in episodic MDPs.
@@ Appendix A Related Work
Reinforcement learning, with a wide range of applications in gaming AIs (Bradley Knox & Stone, 2008; MacGlashan et al., 2017; Warnell et al., 2018) , recommendation systems (Yang et al., 2024; Zeng et al., 2016; Kohli et al., 2013) , autonomous driving (Wei et al., 2024; Schwarting et al., 2018; Kiran et al., 2022) , and large language model (LLM) training (Ziegler et al., 2019), has achieved tremendous success in the past decade. A typical reinforcement learning problem involves an agent and an environment, where at each step, the agent observes the state, takes a certain action, and then receives a reward signal. The state of the environment then transits to another state, and this process continues. In this section, we review works on RL that is relevant to our paper.
@@
Adaptive Stopping Design in RL: in best arm identification, usually, the key to delivering a promising and close to optimal sample complexity performance is to design the stopping criteria which stops the sampling rule as soon as the best action is identifiable. Before “model-based” methods such as TAS (Garivier & Kaufmann, 2016) which uses complicated Chernoff statistics to control the stopping time, there were many “confidence-based” algorithms (Kalyanakrishnan et al., 2012; Kaufmann & Kalyanakrishnan, 2013; Gabillon et al., 2012) focused on constructing high-probability confidence intervals. The stopping time is then when one confidence interval is disjoint from and greater than all the rest, which inspires our design. A similar designing principle is also adopted for bandit algorithms with multiple objectives, e.g., regret minimization while identifying the optimal action (Degenne et al., 2019; Zhang & Ying, 2023). Within this approach, the algorithms can identify the exact optimal policy, while also adapts the sample complexity to the structure of the problem, involving constants such as the value function gap in the theoretical performance bounds.
@@
Reinforcement Learning from Human Feedback: In recent years, the development of LLMs (Wu et al., 2021; Nakano et al., 2021; Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020) has motivated the study of RLHF (Christiano et al., 2017; Kaufmann et al., 2023), which is a generalization of dueling bandits in the MDP setting. Xu et al. (2020) studies the RLHF problem in standard tabular MDPs by reducing the RL problem into bandits of each state. Novoseller et al. (2020) generalizes posterior sampling to preference-based RL and Saha et al. (2023) studies RLHF under trajectory feature and preferences. For MDPs with function approximation, most existing work (Saha et al., 2023; Zhan et al., 2023b; a; Chen et al., 2022; Kong & Yang, 2022) are based on the Bradley-Terry-Luce model and MLE to characterize the human preferences and transform it into rewards. Specifically, in (Wang et al., 2023), the authors build a preference to reward interface that uses confidence balls to decide whether to query human feedback. Zhu et al. (2023) generalizes reward inference from pair-wise comparison to multiple preference rankings. (Kausik et al., 2024) studies RLHF with partially observed rewards and states. All aforementioned RLHF algorithms are value-based, and Du et al. (2024) studies policy gradient with human feedback. To the best of our knowledge, almost all works in literature adopt the pipeline that infers the reward model and then performs classic RL algorithms. Attempts to bypass reward modeling have also been studied. In (Rafailov et al., 2023), the authors proposed the DPO method which optimizes the policy directly from the preference over trajectories. However, DPO is purely empirical and still assumes the Bradley-Terry model to implicitly infer the reward. Moreover, DPO requires the knowledge of a reference policy and thus is only suitable in the fine-tuning phase. Attempts to understand the theoretical performance of DPO and its generalizations are made in (Gheshlaghi Azar et al., 2024), but the authors showed the existence of optima of the loss function, which does not provide any policy optimality and sample complexity guarantees.
@@ C.1 $\mathsf{BSAD}$ for Discounted MDPs
Assume the Condorcet winner exists for all state $s$ and trajectory length $h\leq H$. We let $\overline{\Delta}_{h}^{M}(s,a)$ to be the probability gap for action $a$ and trajectories of length $H-h$ compared to the Condorcet winner of that state $s$, and $p^{\pi}_{h}(s|s^{\prime})$ is the visitation probability of $s$ after $h$ steps starting from state $s^{\prime}$ with policy $\pi$. Both definitions are analog to the definitions in episodic MDPs.
@@
$K=\tilde{\mathcal{O}}\left(\frac{SA^{3}M\log(\frac{1}{\delta})\log^{3}(\frac{1}{\varepsilon})}{(1-\gamma)^{3}\min_{h,s,a}\overline{\Delta}_{h}^{M}(s,a)^{2}\max_{\pi}\min_{s^{\prime}}p^{\pi}_{h}(s|s^{\prime})^{2}}\right).$